Following schedule gives marks obtained by students in statistics in a class. Find out mean marks using short cut method.

Mid-values

5

10

15

20

25

30

No. of students

5

7

9

10

8

6



1
Expert's answer
2021-07-01T05:04:27-0400

Fi is No of students, Xi is mid-point, a = 20, h = 5, di = (Xi-a)/h


Fi Xi (Xi - a) di fi×di

5 5 -15 -3 -15

7 10 -10 -2 -14

9 15 -5 -1 -9

10 20 0 0 0

8 25 5 1 8

6 30 10 2 12

Sum Fi: 45 -18


Mean = 20 + 5×(-18)/45 = 18.